OpenStudy (anonymous):

The original value of a car is $18,000, and it depreciates by 15% each year. what is the value of the car after three years.

OpenStudy (anonymous):

is it 11054.25?

OpenStudy (anonymous):

18,000x.15 = 2700 18000-2700=15300 15300x.15=2295 15300-2295=13005 13005x.15=1950.75 13005-1950.75=11054.25?

OpenStudy (anonymous):

does this look right?????

OpenStudy (anonymous):

i looked at it, took the quiz, and i was right :)

OpenStudy (anonymous):

oops sorry u r actually multiplying by 0.15

OpenStudy (anonymous):

o then what u got should be the right answer

OpenStudy (anonymous):

okay, it's always good to have someone double checking you though